OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). 2x + 3y = –17 5x + 2y = –4 A. (3, –1) B. (–2, 2) C. (–1, 3) D. (2, –7)

OpenStudy (anonymous):

if you can add the equations and one variable will be as big in the first as it will be negative in the second, then that variable gets eliminated

OpenStudy (imstuck):

Let's multiply the first equation by -2 and the second equation by 3. Let see why\[-2(2x+3y=-17)=-4x-6y=34\]\[3(5x+2y=-4)=15x+6y=-12\]The y's will cancel because wee have of them then we are taking 6 of them away. Now we only have equations with x's in them. Let's solve for x. First, here's what we are left with: 11x = 22. Solve for x. When you find x, sub that in for an x in one of the equations you were given at the start and solve for y. If you have problems with this, let us know.